You sell widgets for $10 and have 500 customers. For every .25 cents you raise the price you will lose 2 customers. At what selling price would you have the most revenue and what would that revenue and selling price be ? Also how many customers would you have ?
Answer by stanbon(75887)

You sell widgets for $10 and have 500 customers. For every .25 cents you raise the price you will lose 2 customers. At what selling price would you have the most revenue and what would that revenue and selling price be ? Also how many customers would you have ?
-------
Revenue = (# of units sold)(price per unit)
-----
Equation:
R = (500-2x)(10+0.25x)
R = 5000+125x -20x -0.5x^2
----
R = (-1/2)x^2 +105x + 5000
-----
Vertex:: (105,10512.50)
----------
At what selling price would you have the most revenue?
Ans: $10.00+0.25*105 = $36.25
--------------------
and what would that revenue be?
Ans: $10512.50
---------------------
Also how many customers would you have ?
Ans: 500 - 2*105 = 290
--------------
Cheers,
Stan H.
--------------